Can you please help me what is 4/5 of a pound of peanuts when each pound is $3.00
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

1 pound : 3 dollars


(4/5)*1 pound : (4/5)*3 dollars multiply both sides by 4/5


4/5 pound : 12/5 dollars


4/5 pound : 2.40 dollars

Answer by stanbon(75887)

Can you please help me? What is the value of 4/5 of a pound of peanuts when each pound is $3.00
----------
Ans:: (4/5)3.00 = 4*0.60 = $2.40
==================
Cheers,
Stan H.